OpenStudy (anonymous):

Angle x has a supplementary angle measuring 37 degrees. What is the measure of angle x? Answer: a. 23 degrees b. 37 degrees c. 53 degrees d. 63 degrees e. 143 degrees

OpenStudy (ajprincess):

Sum of supplementary angles=180 x+37=180 Can u fin x @skittles223?

OpenStudy (mayankdevnani):

x+37=180 x=180-37=143

OpenStudy (mayankdevnani):

ok @skittles223